26millecai:

Part B: If (6 and little 0)x = 1, what are the possible values of x? Explain your answer. (5 points)

Astrid1:

\[(a^n)^m=^n*^m\] \[(5^0)^x=1\] \[5^0=1\] Therefore, X can be any value.

26millecai:

thank you!